OpenStudy (anonymous):

If the points (8,1) , (k,-4) and (2,-5) are collinear the k =

OpenStudy (kc_kennylau):

\[\Large m_{AB}=m_{BC}\]

OpenStudy (anonymous):

is m the slope

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

what is AB

OpenStudy (kc_kennylau):

A, B, C are the points respectively

OpenStudy (anonymous):

how to solve it then

OpenStudy (anonymous):

can u just setup the equation so i can solve it i am not ablee to figure out the quation

OpenStudy (kc_kennylau):

\[\frac{-4-1}{k-8}=\frac{-5-(-4)}{2-k}\]

OpenStudy (anonymous):

oh i get thank u!

OpenStudy (anonymous):

i got it

OpenStudy (kc_kennylau):

no problem